# From Evidence to Verdict: An Agent-Based Forensic Framework for AI-Generated Image Detection

### v1

@@ Results
Robustness Analysis
To assess the robustness of our multi-agent detection framework under realistic conditions, we evaluate its performance on perturbed versions of the dataset. The goal of this analysis is to determine whether the framework can maintain consistent accuracy when the input images undergo quality degradation. We apply a series of standard image degradation techniques, including: Blurring. We apply Gaussian blur with a fixed radius of 2, simulating defocus or motion blur; Sharpening. We enhance image sharpness using a sharpening factor of 2.0, which may alter local edge distributions; Gaussian Noise. We add Gaussian noise with mean 0 and variance 2 to simulate sensor-level or environmental noise. These image transformations represent common distortions encountered in real-world image acquisition or compression pipelines. For each perturbation type, we measure the detection performance on the perturbed images and compare the results against the best baseline, GPT-4o.
Table 5 presents the comprehensive performance comparison between our AIFo framework and the GPT-4o baseline across three types of image perturbations: Blurring (Blu), Gaussian noise (Noi), and sharpening (Sha). The results demonstrate that our framework maintains superior performance under these robustness testing conditions.
Table 5: Performance comparison of AIFo vs GPT-4o under different image perturbations on the overall dataset. AIFo superior results are highlighted in bold.
@@
Our AIFo framework consistently outperforms GPT-4o across all perturbation types and most evaluation metrics. Under blurring conditions, both GPT-4o and AIFo experience noticeable performance degradation, as visual analysis tools are inherently sensitive to loss of a great amount of image pixel features. Nevertheless, AIFo still achieves an accuracy of 0.9047, outperforming GPT-4o’s 0.8818. In contrast, under noisy and sharpened conditions, AIFo demonstrates remarkable robustness: its accuracy remains at 0.9710 and 0.9670 respectively, with only minimal decline compared to the clean setting, and consistently surpasses GPT-4o. This robustness can be attributed to the fact that, apart from pre-trained classifiers, the majority of tools integrated in our framework are largely insensitive to the perturbations, thereby stabilizing overall performance.
Evasive Attack Analysis
@@
Reverse-search manipulation 0.8971 0.8690 0.9353 0.9010
@@
Ablation Study
To investigate the individual contributions of different tool categories within our multi-agent framework, we conduct ablation experiments to quantify the impact of each of the four tool types on overall detection performance. To assess the importance of each category, we adopt a leave-one-out strategy: in each ablation run, we disable one tool category while keeping the others unchanged, and evaluate the resulting performance on the full benchmark dataset. Meanwhile, we temporarily disable the debate mechanism to better isolate and focus on the contributions of the tools themselves. By observing the performance drop associated with the removal of each tool category, we can quantify its relative contribution to the framework’s overall effectiveness. This analysis provides insight into which components are most essential for robust and reliable detection, and helps guide future efforts in optimizing or simplifying the system without substantial loss in performance.
Figure 4 presents the performance degradation observed when each tool category is disabled. The results clearly demonstrate that the vision-language model (VLM) tool contributes the most critical evidence to the overall detection pipeline. Removing the VLM tool causes the largest performance degradation across all metrics, with accuracy dropping below 0.85, and recall falling below 0.70. This indicates that direct image analysis from state-of-the-art VLMs remains the main source of evidence for the reasoning agent.
Disabling any of the metadata extraction, reverse image search, or pre-trained classifier modules leads to a slight decrease in both accuracy and F1 score. This indicates that the evidence provided by these tools offers complementary information that enhances the capability of the reasoning agent. Notably, even without these modules, our framework still surpasses the GPT-4o baseline, confirming that these tools contribute complementary signals that strengthen the robustness of the overall reasoning process.
Overall, the ablation study validates the necessity of a diverse Toolbox for reliable detection. These findings highlight that performance gains in our multi-agent framework arise not from a single dominant component, but from the effective integration of cross-source and complementary evidence sources.
@@
Our evaluation demonstrates the advantages of the proposed multi-agent framework: it consistently achieves the highest detection performance on five benchmark datasets in the lab and in-the-wild images from six online platforms, covering AI-generated content produced by at least 20 models. The framework also shows stronger robustness than the best baseline, GPT-4o, against three common perturbations. We find that our framework can dynamically weigh the confidence of evidence gathered from each tool: when metadata and reverse search tools provide valid evidence, it tends to prioritize them; otherwise, it falls back on VLM analysis and traditional classifiers. The debate module further addresses conflicting evidence and prevents the framework from over-relying on a single source of evidence. Altogether, the framework demonstrates greater effectiveness and robustness compared to single-model approaches for AI-generated image detection.
@@
Figure 5: Overview of the memory-augmented reasoning module.
Our current designed detectors operate as stateless processes, analyzing each image independently without leveraging historical context or learning from past decisions. However, forensic analysis is inherently iterative and cumulative, where insights from previous cases can inform future judgments [25]. To address this limitation, we introduce a memory-augmented reasoning module that enables the framework to maintain a knowledge base of past cases and decisions. As illustrated in Figure 5, the module captures, indexes, and retrieves relevant historical context to inform current detection decisions. The designed knowledge base systematically collects both successful and failed detection cases, storing image embedding vectors together with their associated classification results, collected evidence, and reasoning processes. For failed cases, it additionally stores a reflective analysis that captures the causes of misclassification. When analyzing a new image, the system performs similarity-based retrieval over this knowledge base to identify relevant historical cases.
We hypothesize that the introduction of similar historical cases enables the agent to learn from experience and potentially adjust the weights of evidence from different forensic tools based on past performance patterns. Therefore, the enhanced system first queries its accumulated experience to identify potentially relevant historical cases rather than approaching each image as an isolated classification problem.
Implementation Details
Knowledge Base Dataset. To avoid data leakage, we construct the knowledge base image repository using a completely separate set of images from those used in the main benchmark dataset described in  Section 4.1. Following the same collection and sampling methodology as the benchmark dataset, we curate 600 new images for the knowledge base, containing 300 AI images and 300 real images.
Building Knowledge Base. Our knowledge base system is built on a hybrid architecture that combines vector similarity search for content retrieval with structured metadata storage for case management and analysis. The implementation leverages CLIP [42] embeddings to capture high-level semantic and visual features that enable effective similarity matching across images. Each image processed by the framework undergoes feature extraction using the CLIP-ViT-B/32 model, generating 512-dimensional dense vector representations. These embeddings serve as the primary indexing mechanism, enabling efficient similarity-based retrieval of historically relevant cases. The knowledge base maintains separate indices for successful and failed detection cases. For successful cases, the system records the complete set of evidence used in decision-making and the full analytical process. For failed cases, the system records all the above information and additionally stores a structured reflection, providing deeper insight into the causes of misclassification. The reflection is generated by GPT-4.1 to examine the complete analytical pathway and identify potential failure points. This process produces insights covering evidence misinterpretation, tool reliability assessment, and reasoning inconsistencies. These generated reflections create a systematic learning mechanism that helps the system avoid repeating similar analytical errors. Finally, the knowledge base module is integrated seamlessly as one extra forensic tool. During the evidence collection phase, the module is activated.
Inference. Upon receiving a new image, the system computes its CLIP embedding and performs a similarity search over the knowledge base indices to retrieve the top-$k$ most relevant historical cases, with defaults of $k=1$ and a similarity threshold of $0.85$ to ensure only sufficiently close matches are returned. The retrieved cases are then passed to the reasoning agent as additional context, allowing it to leverage past experiences to inform the current analysis.
@@
We did not incorporate the memory module in the main evaluation, as it requires ground-truth labels during the experience accumulation phase. Allowing such label-dependent learning at test time would introduce unfair advantages to our framework, compared to baselines that do not access these labels. Instead, to highlight the potential of the memory module, we directly focus on failure cases from the main evaluation. These are the scenarios where the knowledge base and additional experience are most needed.
To more directly assess its effectiveness, we conduct a targeted evaluation focusing on failure cases from the original framework. Specifically, we identify 50 misclassified images from the benchmark set for which the knowledge base contained semantically similar counterparts. We then re-evaluated these images with the knowledge base module enabled, examining whether the additional contextual evidence could help correct the prior misjudgments. Table 7 summarizes the results by different image types and settings, showing the number of errors successfully corrected after incorporating knowledge base evidence. The results demonstrate that over 40% of error cases are successfully corrected when the memory module is employed. Furthermore, the example illustrated in Figure 6 highlights how retrieved similar cases influenced the reasoning process. When presented with similar past cases, the agents internally reweighted the credibility of different forensic tools, reinterpreted conflicting evidence, and ultimately arrived at the correct conclusion. These observations suggest that the memory-augmented reasoning offers substantial benefits in failure recovery and demonstrates strong potential for enhancing adaptive reasoning in forensic scenarios.
Table 7: Number of errors before and after incorporating similar case history. FP: real images misclassified as AI images; FN: AI images misclassified as real images.
@@
Figure 6: Example of memory-augmented reasoning showing how historical cases influence the agent’s decision-making process and evidence weighting.
@@
Adversarial Metadata Manipulation. Our framework lies in its reliance on image EXIF metadata as a key source of forensic evidence. However, adversaries could potentially manipulate image metadata to mislead the detection system. For instance, attackers could inject fake EXIF data mimicking legitimate camera parameters into AI-generated images. Such metadata spoofing attacks could compromise the reliability of our metadata extraction tool, which currently shows high reliability rates in our evaluation.
Related Work
Fake Image Detection
Early approaches to AI-generated image detection mainly rely on training image classifiers using machine learning, often using datasets produced by specific generative models. Over time, research has shifted towards more generalizable and robust detection strategies, including leveraging the visual capabilities of large multimodal models (LMMs) such as CLIP [42]. CNNSpot [53] is one of the first works to propose a “universal” detector capable of distinguishing real images from CNN-generated ones without being dependent on a particular network architecture or training dataset. De-Fake [46] extends the detection pipeline by incorporating both the image content and generated prompts using CLIP to train a hybrid classifier. DIRE [54] and ZeroFake [47] exploit intrinsic differences between real and fake images revealed during the diffusion model reconstruction process to build detection models with improved generalization. PatchCraft [59] focuses on local texture patches, identifying subtle artifacts left by generative models in fine-grained regions.
Recent works utilize VLMs as fake image detectors, such as AntifakePrompt [18], the method proposed by Jia et al. [29], and the approach by Ji et al. [28]. AIGI-Holmes [60] proposes a complete framework to train a VLM for explainable and generalizable detection, aiming to produce human-verifiable justifications. The work by Yu et al. [57] develops a framework to enhance generalization and explainability by using a knowledge-guided detector and a forgery-aware prompt learner. FakeBench [33] and DFBench [51] introduced a large-scale benchmark to rigorously test the detection performance of LMMs against a wide range of modern generative models.
Despite these progresses, existing methods exhibit inherent limitations. These approaches predominantly rely on internal, pixel-level visual features, while overlooking complementary external sources of evidence that can be critical for reliable forensic analysis. In contrast to standalone detector architectures, our work adopts an agent-based framework that emulates the investigative workflow of human forensic experts. The proposed AI agent leverages a diverse set of external tools to systematically collect, integrate, and reason over different evidence, which allows for a more robust, context-aware, and explainable framework for AI-generated image detection.
LLM-Based Multi-Agent Frameworks
Recent advances in large language models (LLMs) have catalyzed the development of agentic frameworks that simulate complex human-like workflows by coordinating multiple specialized agents. ReAct [56] introduces a framework that switches between reasoning and acting within language models. Other works such as MetaGPT [24] simulate various roles in a software company and build a multi-agent software development framework. In the security domain, several studies have developed multi-agent frameworks tailored to diverse security-related downstream tasks, including adversarial defense [58], harmful content detection [36], bias identification in generative models [52], fake news verification [32], and comparative evaluation of misinformation detection strategies [27]. However, existing approaches do not specifically target zero-shot detection tasks nor address the unique challenges of AI-generated image detection. To the best of our knowledge, our work is the first to develop a training-free, zero-shot multi-agent framework specifically designed for AI-generated image detection. Combining the strengths of existing image detection methods and LLM agents, we present a novel and practical solution for addressing the evolving challenges in AI-generated image forensics.
@@ Conclusion
In this work, we introduce AIFo, a novel multi-agent framework that advances AI-generated image detection by emulating human forensic reasoning. Unlike conventional methods, AIFo integrates diverse forensic tools and uses multi-agent collaboration to synthesize evidence. Our evaluation on a comprehensive benchmark spanning both lab and real-world settings shows that AIFo achieves 97.05% accuracy, substantially outperforming traditional classifiers and state-of-the-art VLMs like GPT-4o.
The key contributions include a training-free, agent-based paradigm that ensures generalizability through procedural reasoning, a finding that multi-source evidence integration and structured debate mechanisms significantly enhance both accuracy and interpretability, and the establishment of a comprehensive benchmark that enables evaluation under real-world conditions. While our framework represents a significant advancement in AI-generated image detection, it also faces some limitations including computational overhead and dependency on external services.
As generative AI technologies continue to evolve rapidly, the need for robust, interpretable, and adaptable detection systems becomes increasingly critical. Our agent-based approach offers a sustainable foundation for addressing these challenges, providing a framework that can generalize alongside advancing generative models while maintaining the transparency and reliability essential for real-world deployment in security-critical applications.
@@ Appendix A Prompts Used in Our Framework
- An AI-generated image refers to images that is fully or partially generated by AI models.
@@
- An AI-generated image refers to images that is fully or partially generated by AI models.
@@
- An AI-generated image refers to images that is fully or partially generated by AI models.
@@ Appendix B Metadata Analysis Tool Key Fields
Appendix C Detailed Accuracy performance
Table 17: Detailed Accuracy performance of different methods on each image sources.
In-the-Lab In-the-Wild
Real Images AI Images Real Images AI Images
CNNSpot [53] 0.9980 0.9980 1.0000 0.0475 0.1857 1.0000 1.0000 0.9980 0.0000 0.0020 0.0000
PatchCraft [59] 0.9900 0.9140 0.7680 0.8525 0.5986 0.8600 0.9040 0.5960 0.0140 0.2360 0.3360
DE-FAKE [46] 0.8980 0.7280 0.3480 0.7225 0.6443 0.7180 0.5700 0.4920 0.9840 0.9380 0.8360
GPT-4.1 [8] 1.0000 0.9920 0.9880 0.8350 0.9086 0.9980 0.9860 0.9920 0.7840 0.9900 0.9500
GPT-4o [39] 1.0000 0.9860 0.9960 0.8850 0.9471 0.9980 0.9820 0.9940 0.7520 0.9900 0.9420
@@ Appendix D AI Model Sources
GenImage [61] BigGAN, GLIDE, VQDM, ADM, Midjourney, Wukong, and Stable Diffusion (v1.4, v1.5).
FakeBench [33] ProGAN, StyleGANs, CogView2, FuseDream, VQDM, GLIDE, Midjourney, Stable Diffusion, DALL·E 2, and DALL·E 3.
@@
Lexica [9] Lexica Aperture Series (v3.5, v4, v5, Max).
Nightcafe [10] DALL·E 2, DALL·E 3, Stable Diffusion, and various other community fine-tuned models.
Civitai [5] A vast collection of community fine-tuned models, predominantly based on Stable Diffusion (including SDXL variants) series.

### v2

@@ Results
Table 4: AIFo vs. GPT-4o performance under image perturbations. AIFo ’s superior results are in bold.
@@
Robustness Analysis
To evaluate robustness of our framework under realistic degradations, we test its performance on perturbed versions of the dataset. Three common distortions are applied: Gaussian blur (radius = 2) to simulate defocus, Gaussian noise (mean = 0, variance = 2) to mimic sensor noise, and sharpening (factor = 2.0) to simulate edge distributions. These perturbations reflect typical degradations in real-world imaging and compression pipelines and we compare the performance against the best baseline, GPT-4o.
As shown in Table 4, both AIFo and GPT-4o experience performance degradation under blur due to the loss of fine-grained visual details. However, AIFo consistently maintains higher accuracy across all perturbation types. We observe that perturbations increase disagreement between visual evidence and other evidence sources, thereby triggering debate mechanism more frequently. This allows the framework to rely less on degraded visual cues and instead emphasize complementary non-visual evidence when available. Under noise and sharpening, where visual distortions are less severe, AIFo exhibits only minor performance drops relative to the clean setting. This behavior suggests that procedural reasoning provides a mechanism for adapting to changes in evidence reliability. Overall, these results indicate that robustness in AIFo emerges from its ability to mediate heterogeneous evidence under varying degrees of visual uncertainty.
@@
Reverse search manipulation 0.8971 0.8690 0.9353 0.9010
@@
This evaluation demonstrates that procedural reasoning offers distinct advantages for AI-generated image detection, particularly in scenarios where evidence is incomplete, ambiguous, or conflicting. By explicitly assessing evidence sufficiency and resolving conflicts through structured reasoning, the framework is able to revise misleading initial judgments and produce more stable and interpretable decisions. At the same time, these benefits come with increased computational cost, underscoring a fundamental trade-off between efficiency and decision robustness. Our results suggest that procedural, agent-based reasoning is most appropriate for forensic and high-stakes applications, where reliability, and auditability are prioritized over real-time inference. Together, these findings support that AI-generated image detection is better viewed as an evidence-driven decision process rather than a purely predictive task, and highlight procedural reasoning as a promising direction for addressing the limitations of existing detectors.
@@
The core procedural reasoning framework operates in a stateless manner, analyzing each image independently without access to prior cases. While this design ensures fairness and prevents information leakage in benchmark evaluation, human forensic reasoning is often cumulative, drawing on past experiences to inform judgments in difficult or ambiguous cases [HRMP25]. Motivated by this observation, we explore a memory-augmented reasoning module as an optional component that provides contextual reference to past cases. The memory module maintains a knowledge base of previously analyzed cases, including their visual embeddings, collected evidence, final decisions, and reasoning traces. For failure cases, it additionally stores structured reflections generated by LLM that describe potential causes of error, such as unreliable evidence sources or misinterpretation of visual cues. During inference, the memory module retrieves semantically similar cases using CLIP-based embedding similarity. These retrieved cases are passed to the reasoning agent as supplementary context, allowing it to reflect on similarities and differences between the current input and past cases. The agent may choose to incorporate this information when reassessing evidence reliability or resolving conflicts, but is not required to follow past decisions. This design reduces the risk of systematic bias or error propagation, as memory serves as a reference instead of a prescriptive rule.
To avoid data leakage, the knowledge base is constructed from a separate set of 600 images (300 real, 300 AI-generated) distinct from the benchmark dataset. Each image is embedded using CLIP-ViT-B/32 [RKHRGASAMCKS21], and indexed for similarity retrieval. The memory module is integrated as an additional forensic tool within the framework and activated during evidence collection. When analyzing a new image, it retrieves the top-$k$ most relevant past cases ($k=1$ by default), which are passed to the reasoning agent as additional context.
@@
We do not include the memory-augmented module in the main evaluation, as constructing the memory knowledge base requires access to ground-truth labels during the experience accumulation phase. Including such label dependent information at test time would introduce an unfair advantage compared to baseline methods. Instead, we evaluate the memory module through a controlled failure recovery analysis, which reflects its intended use as an additional mechanism for difficult cases. Specifically, we analyze 50 misclassified images from the main benchmark for which semantically similar cases exist in the knowledge base. These cases represent scenarios where the stateless reasoning fails and additional contextual reference may be beneficial. After enabling memory-augmented reasoning, over 40% of these previously misclassified samples are corrected (Table 6). Qualitative examples in Figure 5 illustrate how retrieved past cases support reflective reasoning. These observations suggest that the memory-augmented reasoning offers substantial benefits in failure recovery and strong potential for enhancing adaptive reasoning.
Table 6: Number of errors before and after incorporating similar case history. FP: False Positive Cases; FN: False Negative Cases.
@@
Figure 5: Example of memory-augmented reasoning illustrating the impact of similar historical cases on decision-making.
@@
Adversarial Metadata Manipulation. The framework’s reliance on image EXIF metadata as a key source of forensic evidence is a limitation. However, adversaries could potentially manipulate image metadata to mislead the detection system. For instance, attackers could inject fake EXIF data mimicking legitimate camera parameters into AI-generated images. Such metadata spoofing attacks could compromise the reliability of our metadata extraction tool, which currently shows high reliability rates in our evaluation.
@@ Conclusion
This work presents AIFo, a multi-agent procedural reasoning framework that takes AI-generated image detection as an evidence-driven forensic decision process. Instead of proposing yet another detector, we show that explicitly modeling how forensic evidence is collected, evaluated, and reconciled leads to more reliable and interpretable decisions. Extensive experiments demonstrate that integrating heterogeneous evidence within a structured reasoning pipeline enables effective conflict resolution and decision revision. Our results suggest that robustness in AI-generated image detection is better achieved through multi-stage reasoning over evidence than through increasingly complex detection models alone. We believe that procedural, agent-based reasoning provides a promising foundation for developing forensic systems that remain effective as generative models continue to evolve.
Limitation. The proposed framework has higher computational cost than single-model baselines due to multi-stage reasoning and external tool invocation. While our analysis shows that conflict resolution reasoning is only required for a subset of inputs, further optimization will be important for large-scale deployment. In addition, parts of the reasoning process, including the assessment of evidence sufficiency and consistency, rely on qualitative judgments produced by LLMs and may introduce variability across models or prompts. Exploring hybrid approaches that combine procedural reasoning with more formal uncertainty estimation and robustness guarantees is a promising avenue for future research.
@@ Appendix A Prompts Used in Our Framework
- An AI-generated image refers to images that are fully or partially generated by AI models.
@@
- An AI-generated image refers to images that are fully or partially generated by AI models.
@@
- An AI-generated image refers to images that are fully or partially generated by AI models.
@@ Appendix B Metadata Analysis Tool Key Fields
Appendix C List of Selected Models from Hugging Face
We select the top five most downloaded classification models for AI-generated image detection available on Hugging Face:
  • haywoodsloan/ai-image-detector-deploy
  • Organika/sdxl-detector
  • legekka/AI-Anime-Image-Detector-ViT
  • Smogy/SMOGY-Ai-images-detector
  • NYUAD-ComNets/NYUAD_AI-generated_images_detector
All models are publicly available on the Hugging Face Hub.
@@ Appendix D AI Model Sources
GenImage [ZCYHLLTHHW23] BigGAN, GLIDE, VQDM, ADM, Midjourney, Wukong, and Stable Diffusion (v1.4, v1.5).
FakeBench [LLWLWRL24] ProGAN, StyleGANs, CogView2, FuseDream, VQDM, GLIDE, Midjourney, Stable Diffusion, DALL·E 2, and DALL·E 3.
@@
Lexica Lexica Aperture Series (v3.5, v4, v5, Max).
Nightcafe DALL·E 2, DALL·E 3, Stable Diffusion, and various other community fine-tuned models.
Civitai A vast collection of community fine-tuned models, predominantly based on Stable Diffusion (including SDXL variants) series.
Appendix E Detailed Accuracy Performance
Table 17: Detailed Accuracy performance of different methods on each image source.
In-the-Lab In-the-Wild
Real Images AI Images Real Images AI Images
CNNSpot [WWZOE20] 0.9980 0.9980 1.0000 0.0475 0.1857 1.0000 1.0000 0.9980 0.0000 0.0020 0.0000
PatchCraft [ZXLQZ24] 0.9900 0.9140 0.7680 0.8525 0.5986 0.8600 0.9040 0.5960 0.0140 0.2360 0.3360
DE-FAKE [SLYZ22] 0.8980 0.7280 0.3480 0.7225 0.6443 0.7180 0.5700 0.4920 0.9840 0.9380 0.8360
GPT-4o [GPT4o] 1.0000 0.9860 0.9960 0.8850 0.9471 0.9980 0.9820 0.9940 0.7520 0.9900 0.9420